Bonita Industries sells its product for $7100 per unit. Variable costs per unit are: manufacturing, $3000, and selling and administrative, $100. Fixed costs are: $18000 manufacturing overhead, and $24000 selling and administrative. There was no beginning inventory at 1/1/15. Production was 20 units per year in 2015–2017. Sales were 20 units in 2015, 16 units in 2016, and 24 units in 2017. Income under absorption costing for 2016 is
Answer #1
Fixed manufacturing overhead per unit 900 =18000/20
Unit product cost 3900 =3000+900
Sales revenue 113600 =16*7100
Less: Cost of goods sold 62400 =16*3900
Gross margin 51200
Less: Selling and administrative expenses 25600 =24000+(16*100)
Income under absorption costing for 2016 25600